# Where MLLMs Attend and What They Rely On: Explaining Autoregressive Token Generation

### v2

@@ 4.1 Experimental Setup
Datasets. We evaluate across three representative tasks: MS COCO Caption (Lin et al., 2014; Chen et al., 2015) for image captioning, MMVP (Tong et al., 2024) for visual question answering (VQA), and RePOPE (Neuhaus & Hein, 2025) for object hallucination assessment.
Baselines. We compare Eagle against state-of-the-art attribution methods for MLLMs, including gradient-based approaches (LLaVA-CAM (Zhang et al., 2025b) and IGOS++ adaptation (Xing et al., 2025)) and the activation-based method TAM (Li et al., 2025a). Note that TAM is restricted to attributing a single token at a time and cannot handle token combinations.
Models. We validate our approach on three multimodal large language models: LLaVA-1.5-7B (Liu et al., 2024), Qwen2.5-VL (3B and 7B) (Bai et al., 2025), and InternVL 3.5-4B (Wang et al., 2025).
Evaluation Metrics. We consider three categories of attribution metrics: faithfulness, localization, and correction-oriented. (1) Faithfulness metrics evaluate whether explanations align with the model’s decision process. We adopt Insertion (Petsiuk et al., 2018), Deletion (Petsiuk et al., 2018), and Average Highest Score (Chen et al., 2024b), computed as the mean probability over selected tokens. (2) Localization metrics assess whether explanations overlap with ground-truth regions using the Point Game (Zhang et al., 2018), under both box-level and mask-level annotations, where correctness is defined by the maximum attribution point falling inside the bounding box or segmentation mask. (3) Correction-oriented metrics address hallucination evaluation by testing whether attributions reveal regions causing hallucinated outputs. We use Average Minimal Correction Region (AMCR), the average proportion of regions that must be removed to correct hallucinations, and Correction Success Rate under Budget (CSR@10%), the percentage of cases corrected when no more than 10% of regions are removed.
@@
Datasets MLLMs Methods Sentence-level Faithfulness Sensitive Tokens-level Faithfulness GPU Memory ($\downarrow$)
@@
MS COCO (Lin et al., 2014)
(Image caption task)
LLaVA-1.5 7B
(Liu et al., 2024)
LLaVA-CAM (Zhang et al., 2025b) 0.5298 0.5317 0.6031 0.4124 0.4115 0.5783 37.25 GB
iGOS++ (w/ GNC) (Xing et al., 2025) 0.5293 0.5168 0.6004 0.4101 0.3815 0.5731 48.18 GB
Eagle 0.5970 0.4554 0.6259 0.5344 0.2809 0.5993 16.07 GB
Qwen2.5-VL 3B
(Bai et al., 2025)
LLaVA-CAM (Zhang et al., 2025b) 0.4978 0.5562 0.6662 0.3541 0.4497 0.6424 28.99 GB
iGOS++ (w/ GNC) (Xing et al., 2025) 0.5328 0.4891 0.6672 0.4021 0.3273 0.6473 71.62 GB
Eagle 0.6479 0.4345 0.7039 0.5867 0.2710 0.6840 8.75 GB
Qwen2.5-VL 7B
(Bai et al., 2025)
LLaVA-CAM (Zhang et al., 2025b) 0.5605 0.5464 0.7235 0.4467 0.4209 0.7010 47.17 GB
iGOS++ (w/ GNC) (Xing et al., 2025) 0.5603 0.5072 0.7237 0.4400 0.3623 0.6695 96.90 GB
Eagle 0.7006 0.4597 0.7578 0.6337 0.2988 0.7285 17.68 GB
InternVL3.5 4B
(Wang et al., 2025)
LLaVA-CAM (Zhang et al., 2025b) 0.6116 0.6235 0.8032 0.4948 0.5100 0.7764 81.84 GB
iGOS++ (w/ GNC) (Xing et al., 2025) 0.6271 0.5726 0.7999 0.5088 0.4337 0.7715 60.93 GB
Eagle 0.7665 0.4650 0.8335 0.7042 0.3042 0.8051 12.45 GB
MMVP (Tong et al., 2024)
(VQA task)
LLaVA-1.5 7B
(Liu et al., 2024)
LLaVA-CAM (Zhang et al., 2025b) 0.7756 0.7745 0.7980 0.6076 0.6044 0.7275 34.38 GB
iGOS++ (w/ GNC) (Xing et al., 2025) 0.7717 0.7698 0.7965 0.5825 0.5781 0.7236 92.90 GB
Eagle 0.7960 0.7474 0.8086 0.6867 0.5027 0.7507 15.40 GB
Qwen2.5-VL 3B
(Bai et al., 2025)
LLaVA-CAM (Zhang et al., 2025b) 0.7742 0.7770 0.8181 0.5925 0.6006 0.7476 19.17 GB
iGOS++ (w/ GNC) (Xing et al., 2025) 0.7719 0.7613 0.8183 0.5719 0.5356 0.7437 19.79 GB
Eagle 0.8052 0.7338 0.8339 0.6634 0.4935 0.7689 8.76 GB
Qwen2.5-VL 7B
(Bai et al., 2025)
LLaVA-CAM (Zhang et al., 2025b) 0.7505 0.7486 0.8042 0.4974 0.4847 0.7242 37.54 GB
iGOS++ (w/ GNC) (Xing et al., 2025) 0.7394 0.7211 0.8036 0.4505 0.3853 0.7185 32.76 GB
Eagle 0.7824 0.6996 0.8119 0.5901 0.3675 0.7362 17.40 GB
InternVL3.5 4B
(Wang et al., 2025)
LLaVA-CAM (Zhang et al., 2025b) 0.7348 0.7458 0.8325 0.4897 0.5213 0.7575 27.20 GB
iGOS++ (w/ GNC) (Xing et al., 2025) 0.7277 0.7160 0.8302 0.4743 0.4454 0.7535 62.31 GB
Eagle 0.8012 0.6782 0.8471 0.6379 0.4027 0.7762 12.26 GB
@@ 4.2 Faithfulness on Sentence-level Explanations
We begin by evaluating our attribution method on two common MLLM tasks, image captioning and visual question answering (VQA), with the goal of identifying which image regions drive the full content generated by the model. We primarily compare our approach against LLaVA-CAM (Zhang et al., 2025b) and IGOS++ (w/ GNC) (Xing et al., 2025). Table 1 reports results on faithfulness metrics, evaluated in two ways: (1) using the sum of logits over all predicted tokens, and (2) using the sum over sensitive tokens, defined as those whose logits change by more than 0.2 when the entire image is masked.
@@
In addition to higher attribution fidelity, Eagle demonstrates strong efficiency, requiring only 17.68 GB on Qwen2.5-VL 7B compared to 96.90 GB for IGOS++, making it practical for modern MLLMs. Overall, it provides more faithful and resource-efficient explanations than gradient-based baselines. As shown in Fig. 3, LLaVA-CAM often misses key regions and IGOS++ yields redundant maps, while our method highlights critical regions that align closely with visually grounded tokens, producing concise and human-consistent explanations.
@@
Datasets MLLMs Methods Word-level Faithfulness Metrics Localization Metrics GPU Memory ($\downarrow$)
@@
MS COCO (Lin et al., 2014)
(Image caption task)
LLaVA-1.5 7B
(Liu et al., 2024)
LLaVA-CAM (Zhang et al., 2025b) 0.4063 0.4035 0.6053 0.2468 0.1168 36.73 GB
iGOS++ (w/ GNC) (Xing et al., 2025) 0.4093 0.3812 0.6084 0.6623 0.5584 93.12 GB
TAM (Li et al., 2025a) 0.3860 0.4162 0.5988 0.1818 0.1428 16.60 GB
Eagle 0.6395 0.2047 0.7213 0.8052 0.7792 16.31 GB
Qwen2.5-VL 3B
(Bai et al., 2025)
LLaVA-CAM (Zhang et al., 2025b) 0.3417 0.4575 0.7263 0.1045 0.0621 26.01 GB
iGOS++ (w/ GNC) (Xing et al., 2025) 0.4141 0.2901 0.7250 0.5822 0.4967 58.1 GB
TAM (Li et al., 2025a) 0.5130 0.2797 0.7985 0.5294 0.4379 9.56 GB
Eagle 0.7353 0.1628 0.8641 0.8104 0.7745 9.22 GB
Qwen2.5-VL 7B
(Bai et al., 2025)
LLaVA-CAM (Zhang et al., 2025b) 0.4170 0.4771 0.8041 0.2176 0.1428 44.26 GB
iGOS++ (w/ GNC) (Xing et al., 2025) 0.4816 0.3478 0.8080 0.6734 0.5959 82.14 GB
TAM (Li et al., 2025a) 0.5768 0.3167 0.8240 0.5369 0.4060 18.75 GB
Eagle 0.8109 0.2127 0.9194 0.7785 0.7383 18.03 GB
InternVL3.5 4B
(Wang et al., 2025)
LLaVA-CAM (Zhang et al., 2025b) 0.4988 0.5040 0.8588 0.3201 0.2212 81.84 GB
iGOS++ (w/ GNC) (Xing et al., 2025) 0.5192 0.3983 0.8604 0.5775 0.5181 60.06 GB
TAM (Li et al., 2025a) 0.6317 0.3517 0.8712 0.5775 0.4653 14.23 GB
Eagle 0.8623 0.1706 0.9585 0.8052 0.7755 7.61 GB
4.3 Faithfulness and Localization on Word-level Explanations
Next, we evaluate the ability of the proposed attribution method to provide word-level explanations. Specifically, we use samples with object bounding box annotations from the MS COCO dataset to verify whether the objects mentioned in image captions are accurately grounded in the visual input. We also include TAM (Li et al., 2025a) as an additional baseline, since it is particularly effective at explaining object localization.
For localization, our method achieves the best Pointing Game results under both box- and mask-level settings, confirming that predictions are grounded in specific objects. While TAM performs well on stronger models but poorly on LLaVA-1.5, IGOS++ gains from overly redundant maps. In contrast, our method yields sparse yet focused highlights that more accurately localize the objects mentioned in captions (Fig. 4).
@@ 4.3 Faithfulness and Localization on Word-level Explanations
RePOPE (Neuhaus & Hein, 2025)
(Object Hallucination Benchmark)
LLaVA-1.5 7B
(Liu et al., 2024)
LLaVA-CAM (Zhang et al., 2025b) 0.4095 0.4191 0.6596 0.5613 19.70% 37.07 GB
iGOS++ (w/ GNC) (Xing et al., 2025) 0.4232 0.4182 0.6794 0.4770 37.50% 93.88 GB
TAM (Li et al., 2025a) 0.4168 0.4166 0.6705 0.5826 18.75% 16.59 GB
@@
Qwen2.5-VL 3B
(Bai et al., 2025)
LLaVA-CAM (Zhang et al., 2025b) 0.3994 0.3783 0.6992 0.4555 42.21% 27.10 GB
iGOS++ (w/ GNC) (Xing et al., 2025) 0.4056 0.4471 0.7235 0.4461 37.57% 35.16 GB
TAM (Li et al., 2025a) 0.3905 0.4090 0.6900 0.4747 29.75% 9.66 GB
@@
Qwen2.5-VL 7B
(Bai et al., 2025)
LLaVA-CAM (Zhang et al., 2025b) 0.2444 0.2901 0.5898 0.6620 35.37% 45.05 GB
iGOS++ (w/ GNC) (Xing et al., 2025) 0.3017 0.3330 0.7125 0.5357 32.41% 70.86 GB
TAM (Li et al., 2025a) 0.2717 0.3177 0.6792 0.5844 22.45% 18.57 GB
@@
InternVL3.5 4B
(Wang et al., 2025)
LLaVA-CAM (Zhang et al., 2025b) 0.4079 0.3733 0.9296 0.4078 36.43% 87.93 GB
iGOS++ (w/ GNC) (Xing et al., 2025) 0.3651 0.4556 0.9393 0.4299 38.76% 66.26 GB
TAM (Li et al., 2025a) 0.3794 0.4221 0.9115 0.4801 28.57% 14.04 GB
@@
4.4 Interpreting Object Hallucination
We next apply our interpretable algorithm to analyze why MLLMs produce hallucinations. Experiments are conducted on the object hallucination benchmark RePOPE (Neuhaus & Hein, 2025). We focus on samples where the MLLM makes prediction errors, including cases where the model incorrectly answers “no” instead of “yes,” and vice versa. Assuming that hallucinations have already been identified, our objective is to identify which image regions trigger the hallucination and to assess whether blocking these regions can mitigate it. In practice, we attribute the first token of the answer, restricted to the vocabulary IDs ’Yes’ and ’No’. For example, if the model incorrectly outputs ’Yes’, the attribution is computed with respect to ’No’, thereby providing a counterfactual perspective on which regions would support the correct response.

### v4

@@ 4.1 Experimental Setup
Datasets. We evaluate across three representative tasks: MS COCO Caption [23, 11] for image captioning, MMVP [30] for visual question answering (VQA), and RePOPE [27] for object hallucination assessment.
Baselines. We compare Eagle against state-of-the-art attribution methods for MLLMs, including gradient-based approaches (LLaVA-CAM [36] and IGOS++ adaptation [33]) and the activation-based method TAM [17]. Note that TAM is restricted to attributing a single token at a time and cannot handle token combinations.
Models. We validate our approach on three multimodal large language models: LLaVA-1.5-7B [25], Qwen2.5-VL (3B and 7B) [4], and InternVL 3.5-4B [32].
@@
Datasets MLLMs Methods Sentence-level Faithfulness Sensitive Tokens-level Faithfulness GPU Memory ($\downarrow$) Time Consum. ($\downarrow$)
@@
MS COCO [23] (Image caption task) LLaVA-1.5 7B [25] LLaVA-CAM [36] 0.5298 0.5317 0.6031 0.4124 0.4115 0.5783 37.25 GB 15.4
iGOS++ (w/ GNC) [33] 0.5293 0.5168 0.6004 0.4101 0.3815 0.5731 48.18 GB 36.3 s
Eagle 0.5970 0.4554 0.6259 0.5344 0.2809 0.5993 16.07 GB 258.4 s
Qwen2.5-VL 3B [4] LLaVA-CAM [36] 0.4978 0.5562 0.6662 0.3541 0.4497 0.6424 28.99 GB 22.2 s
iGOS++ (w/ GNC) [33] 0.5328 0.4891 0.6672 0.4021 0.3273 0.6473 71.62 GB 46.6 s
Eagle 0.6479 0.4345 0.7039 0.5867 0.2710 0.6840 8.75 GB 418.3 s
Qwen2.5-VL 7B [4] LLaVA-CAM [36] 0.5605 0.5464 0.7235 0.4467 0.4209 0.7010 47.17 GB 21.8 s
iGOS++ (w/ GNC) [33] 0.5603 0.5072 0.7237 0.4400 0.3623 0.6695 96.90 GB 62.7 s
Eagle 0.7006 0.4597 0.7578 0.6337 0.2988 0.7285 17.68 GB 436.2 s
InternVL3.5 4B [32] LLaVA-CAM [36] 0.6116 0.6235 0.8032 0.4948 0.5100 0.7764 81.84 GB 54.3 s
iGOS++ (w/ GNC) [33] 0.6271 0.5726 0.7999 0.5088 0.4337 0.7715 60.93 GB 40.1 s
Eagle 0.7665 0.4650 0.8335 0.7042 0.3042 0.8051 12.45 GB 803.5 s
MMVP [30] (VQA task) LLaVA-1.5 7B [25] LLaVA-CAM [36] 0.7756 0.7745 0.7980 0.6076 0.6044 0.7275 34.38 GB 12.2 s
iGOS++ (w/ GNC) [33] 0.7717 0.7698 0.7965 0.5825 0.5781 0.7236 92.90 GB 34.8 s
Eagle 0.7960 0.7474 0.8086 0.6867 0.5027 0.7507 15.40 GB 252.6 s
Qwen2.5-VL 3B [4] LLaVA-CAM [36] 0.7742 0.7770 0.8181 0.5925 0.6006 0.7476 19.17 GB 16.1 s
iGOS++ (w/ GNC) [33] 0.7719 0.7613 0.8183 0.5719 0.5356 0.7437 19.79 GB 31.8 s
Eagle 0.8052 0.7338 0.8339 0.6634 0.4935 0.7689 8.76 GB 222.7 s
Qwen2.5-VL 7B [4] LLaVA-CAM [36] 0.7505 0.7486 0.8042 0.4974 0.4847 0.7242 37.54 GB 14.5 s
iGOS++ (w/ GNC) [33] 0.7394 0.7211 0.8036 0.4505 0.3853 0.7185 32.76 GB 31.0 s
Eagle 0.7824 0.6996 0.8119 0.5901 0.3675 0.7362 17.40 GB 220.8 s
InternVL3.5 4B [32] LLaVA-CAM [36] 0.7348 0.7458 0.8325 0.4897 0.5213 0.7575 27.20 GB 20.7 s
iGOS++ (w/ GNC) [33] 0.7277 0.7160 0.8302 0.4743 0.4454 0.7535 62.31 GB 34.4 s
Eagle 0.8012 0.6782 0.8471 0.6379 0.4027 0.7762 12.26 GB 176.5 s
Evaluation Metrics. We consider three categories of attribution metrics: faithfulness, localization, and correction-oriented. (1) Faithfulness metrics evaluate whether explanations align with the model’s decision process. We adopt Insertion [29], Deletion [29], and Average Highest Score [10], computed as the mean probability over selected tokens. (2) Localization metrics assess whether explanations overlap with ground-truth regions using the Point Game [34], under both box-level and mask-level annotations, where correctness is defined by the maximum attribution point falling inside the bounding box or segmentation mask. (3) Correction-oriented metrics address hallucination evaluation by testing whether attributions reveal regions causing hallucinated outputs. We use Average Minimal Correction Region (AMCR), the average proportion of regions that must be removed to correct hallucinations, and Correction Success Rate under Budget (CSR@10%), the percentage of cases corrected when no more than 10% of regions are removed.
@@ 4.2 Faithfulness on Sentence-level Explanations
We begin by evaluating our attribution method on two common MLLM tasks, image captioning and visual question answering (VQA), with the goal of identifying which image regions drive the full content generated by the model. We primarily compare our approach against LLaVA-CAM [36] and IGOS++ (w/ GNC) [33]. Table 1 reports results on faithfulness metrics, evaluated in two ways: (1) using the sum of logits over all predicted tokens, and (2) using the sum over sensitive tokens, defined as those whose logits change by more than 0.2 when the entire image is masked.
@@
Datasets MLLMs Methods Word-level Faithfulness Metrics Localization Metrics GPU Memory ($\downarrow$) Time Consum. ($\downarrow$)
@@
MS COCO [23] (Image caption task) LLaVA-1.5 7B [25] LLaVA-CAM [36] 0.4063 0.4035 0.6053 0.2468 0.1168 36.73 GB 17.1 s
iGOS++ (w/ GNC) [33] 0.4093 0.3812 0.6084 0.6623 0.5584 93.12 GB 33.5 s
TAM [17] 0.3860 0.4162 0.5988 0.1818 0.1428 16.60 GB 1.7 s
Eagle 0.6395 0.2047 0.7213 0.8052 0.7792 16.31 GB 283.2 s
Qwen2.5-VL 3B [4] LLaVA-CAM [36] 0.3417 0.4575 0.7263 0.1045 0.0621 26.01 GB 12.8 s
iGOS++ (w/ GNC) [33] 0.4141 0.2901 0.7250 0.5822 0.4967 58.1 GB 49.9 s
TAM [17] 0.5130 0.2797 0.7985 0.5294 0.4379 9.56 GB 4.2 s
Eagle 0.7353 0.1628 0.8641 0.8104 0.7745 9.22 GB 371.4 s
Qwen2.5-VL 7B [4] LLaVA-CAM [36] 0.4170 0.4771 0.8041 0.2176 0.1428 44.26 GB 24.5 s
iGOS++ (w/ GNC) [33] 0.4816 0.3478 0.8080 0.6734 0.5959 82.14 GB 53.1 s
TAM [17] 0.5768 0.3167 0.8240 0.5369 0.4060 18.75 GB 4.9 s
Eagle 0.8109 0.2127 0.9194 0.7785 0.7383 18.03 GB 435.6 s
InternVL3.5 4B [32] LLaVA-CAM [36] 0.4988 0.5040 0.8588 0.3201 0.2212 81.84 GB 41.3 s
iGOS++ (w/ GNC) [33] 0.5192 0.3983 0.8604 0.5775 0.5181 60.06 GB 32.7 s
TAM [17] 0.6317 0.3517 0.8712 0.5775 0.4653 14.23 GB 4.6 s
Eagle 0.8623 0.1706 0.9585 0.8052 0.7755 7.61 GB 377.3 s
In addition to higher attribution fidelity, Eagle demonstrates strong efficiency, requiring only 17.68 GB on Qwen2.5-VL 7B compared to 96.90 GB for IGOS++, making it practical for modern MLLMs. Overall, it provides more faithful and resource-efficient explanations than gradient-based baselines. Although our method is roughly ten times more time-consuming than LLaVA-CAM and I-GOS++, it is the only one that shows promising potential to meet real-world interpretability requirements. As shown in Fig. 3, LLaVA-CAM often misses key regions and IGOS++ yields redundant maps, while our method highlights critical regions that align closely with visually grounded tokens, producing concise and human-consistent explanations.
4.3 Experiments on Word-level Explanations
Next, we evaluate the ability of the proposed attribution method to provide word-level explanations. Specifically, we use samples with object bounding box annotations from the MS COCO dataset to verify whether the objects mentioned in image captions are accurately grounded in the visual input. We also include TAM [17] as an additional baseline, since it is particularly effective at explaining object localization.
@@ 4.3 Experiments on Word-level Explanations
RePOPE [27] (Object Hallucination Benchmark) LLaVA-1.5 7B [25] LLaVA-CAM [36] 0.4095 0.4191 0.6596 0.5613 19.70% 37.07 GB
iGOS++ (w/ GNC) [33] 0.4232 0.4182 0.6794 0.4770 37.50% 93.88 GB
TAM [17] 0.4168 0.4166 0.6705 0.5826 18.75% 16.59 GB
@@
Qwen2.5-VL 3B [4] LLaVA-CAM [36] 0.3994 0.3783 0.6992 0.4555 42.21% 27.10 GB
iGOS++ (w/ GNC) [33] 0.4056 0.4471 0.7235 0.4461 37.57% 35.16 GB
TAM [17] 0.3905 0.4090 0.6900 0.4747 29.75% 9.66 GB
@@
Qwen2.5-VL 7B [4] LLaVA-CAM [36] 0.2444 0.2901 0.5898 0.6620 35.37% 45.05 GB
iGOS++ (w/ GNC) [33] 0.3017 0.3330 0.7125 0.5357 32.41% 70.86 GB
TAM [17] 0.2717 0.3177 0.6792 0.5844 22.45% 18.57 GB
@@
InternVL3.5 4B [32] LLaVA-CAM [36] 0.4079 0.3733 0.9296 0.4078 36.43% 87.93 GB
iGOS++ (w/ GNC) [33] 0.3651 0.4556 0.9393 0.4299 38.76% 66.26 GB
TAM [17] 0.3794 0.4221 0.9115 0.4801 28.57% 14.04 GB
@@
For localization, our method achieves the best Pointing Game results under both box- and mask-level settings, confirming that predictions are grounded in specific objects. While TAM performs well on stronger models but poorly on LLaVA-1.5, IGOS++ gains from overly redundant maps. In contrast, our method yields sparse yet focused highlights that more accurately localize the objects mentioned in captions (Fig. 4). Although TAM [17] is much faster in attribution time, it heavily relies on the MLLM’s activation maps. As a result, while it performs well on Qwen2.5-VL and InternVL 3.5, its effectiveness drops sharply on LLaVA-1.5. In contrast, our method does not depend on model-specific internal activations and therefore remains consistently superior.
4.4 Interpreting Object Hallucination
We next apply our interpretable algorithm to analyze the causes of hallucinations in MLLMs. Experiments are conducted on the object hallucination benchmark RePOPE [27]. Our focus is on samples where the MLLM makes prediction errors, particularly in cases where the model incorrectly answers "no" instead of "yes," and vice versa. The primary aim is not to mitigate hallucinations, but to explain why they occur, specifically, by identifying the image regions that trigger hallucinations. Assuming that hallucinations have already been identified, we aim to pinpoint which image regions are responsible for these errors and to evaluate whether blocking these regions can reduce the hallucinations. Practically, we attribute the first token of the answer, focusing on the vocabulary IDs ’Yes’ and ’No’. For example, if the model incorrectly outputs ’Yes’, the attribution is computed with respect to ’No’, providing a counterfactual perspective on which regions would support the correct response.